OpenStudy (anonymous):

How many solutions does the equation y + 12 = y + 10 + 2 have? Zero One Two Infinitely many

OpenStudy (anonymous):

@AnitaS. @Daniee_Bruhh

OpenStudy (welshfella):

if we add the 10 and the 2 on the right side we get y + 12 = y + 12

OpenStudy (anonymous):

12

OpenStudy (anitas.):

Simplify 10+2 to 12 12=12

OpenStudy (welshfella):

now i can make y = 1 , 2, 5 , 78 - any number i like and the left and right sides of the equation will be equal

OpenStudy (anonymous):

alright so whats the answer

OpenStudy (welshfella):

well i've told you in my last post

OpenStudy (welshfella):

check out the choices

OpenStudy (anonymous):

so its d

OpenStudy (welshfella):

yes infinitely mean solutions

OpenStudy (welshfella):

* many

OpenStudy (anonymous):

thanks so much hears a medal

OpenStudy (anonymous):

can u help me with a couple more

OpenStudy (welshfella):

yes plz repost though